# 1(2), Leninskie gory, Moscow 119991, Russian Federation

### v2

@@ 5 Examples
Figure 1 presents “hard” $\overline{\lambda}_{a}$ (“soft” $\tilde{\lambda}_{a}$ can be recovered using eq. (32)) for leading sectors of some integral families upto the 7-loop order. Polynomials of the first degree as coefficients were sufficient to reduce the original system of equations to an equation for one variable. [^1] Symbolic manipulations were performed using REDUCE (https://reduce-algebra.sourceforge.io/).
$L=5$$c$$e$$e$$c$$e$$e$$b$$b$$a$$b$$b$$c$
$L=6$$c$$e$$e$$e$$c$$e$$e$$e$$b$$b$$a$$b$$b$$c$$c$
$L=7$$c$$e$$e$$e$$e$$c$$e$$e$$e$$e$$b$$b$$a$$b$$b$$c$$c$$c$
$L=6$$e$$e$$e$$e$$e$$e$$c$$c$$c$$c$$b$$b$$a$$b$$b$
$L=7$$e$$e$$e$$e$$e$$e$$e$$e$$c$$c$$c$$c$$c$$b$$b$$a$$b$$b$
Figure 1: $\overline{\lambda}_{a}$ (roots of eqs. (24)) for some vacuum diagrams. Ordinary lines are massless, double line is massive; $a_{L}=L$, $b_{L}=-\frac{L-2}{2}e_{L}$, $c_{L}=\frac{L(L-2)}{2}e_{L}-L(L-1)$;
@@
The results show that in 5-, 6-loop planar examples there is at least one $1/d$ solution, in 7-loop planar and 6-, 7-loop nonplanar cases there are at least two. The $L$-dependence of the results was obtained empirically, but there are some indications of its validity for an arbitrary order, we hope to clarify this point in future publications.
Please note that eq. (16) from Baikov:2005nv (eq. (34) in the notation of this article) at the 7-loop level leads to polynomial equations of the 6th degree for 10 variables, which is a much more difficult task.
@@ 6 Conclusion
In this paper we present a method to construct solutions of IBP relations for Feynman integrals as formal series in the limit $1/d\rightarrow 0$. The idea is to use substitutions of the form (1517) which transform IBP relations into an explicit recursive form. From a technical point of view, it is necessary to calculate the constants $\lambda$ as solutions to a system of bilinear quadratic equations (24). Although this is relatively simple (and therefore can be automated) in the case where one line is massive and the others are massless, more general kinematic configurations may be more difficult to study. In particular, it may happen that the number of $1/d$ solutions constructed will be less than the number of master integrals and, therefore, not enough to calculate all $c_{i}(\underline{n})$ coefficients in eq. (7). However, even in this case they provide some linear relations between $c_{i}(\underline{n})$ that can be used (in combination with other data) for their reconstruction or at least for verification.
Please note that although in this article we are dealing with recurrence relations for vacuum integrals, the substitutions (15, 17) can also be applied to the case with external momenta; we leave this for future research.

### v3

@@ 5 Examples
Let us consider 2-loop example with inverse propagators
$D_{1}=p_{1}^{2},\quad D_{2}=p_{2}^{2},\quad D_{3}=(p_{1}+p_{2})^{2}-1.$
Recurrence relations (5) are
$R_{11}(\underline{n})$ $=$ $1/2(d-2\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}-\overline{\bf I}_{1}^{-}\overline{\bf I}_{3}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{+})f(\underline{n}),$
$R_{12}(\underline{n})$ $=$ $1/2(\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}+\overline{\bf I}_{1}^{-}\overline{\bf I}_{3}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{1}^{+}-\overline{\bf I}_{2}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{1}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{1}^{+}-\overline{\bf I}_{3}^{+})f(\underline{n}),$
$R_{21}(\underline{n})$ $=$ $1/2(\overline{\bf I}_{1}^{-}\overline{\bf I}_{2}^{+}-\overline{\bf I}_{1}^{-}\overline{\bf I}_{3}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{2}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{2}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{2}^{+}-\overline{\bf I}_{3}^{+})f(\underline{n}),$
$R_{22}(\underline{n})$ $=$ $1/2(d+\overline{\bf I}_{1}^{-}\overline{\bf I}_{3}^{+}-2\overline{\bf I}_{2}^{-}\overline{\bf I}_{2}^{+}-\overline{\bf I}_{2}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{+})f(\underline{n}).$ (36)
Let us assume that all indices are “hard”,
[…]
and substitute (remember that $d/2=\alpha^{-2}$):
$f(\underline{n})$ $=$ $\exp(\alpha^{-2}\sum_{a=1}^{3}\overline{\lambda}_{a}\overline{\bf I}_{a}^{-})f^{\prime}(\underline{n}),\quad f^{\prime}(\underline{n})=\alpha^{3-\sum_{a=1}^{3}n_{a}}g(\underline{n}),$ (38)
$f^{\prime}(\underline{n}),\,g(\underline{n})$ $=$ $0\,\,\,\mbox{if any of}\,\,(n_{1},n_{2},n_{3})<1.$ (39)
These substitutions correspond to
$\overline{{\bf I}}_{f,a}^{+}=\alpha^{-1}\overline{{\bf I}}_{g,a}^{+}+\alpha^{-2}\overline{\lambda}_{a},\quad\overline{{\bf I}}_{f,a}^{-}=\alpha\overline{{\bf I}}_{g,a}^{-}.$
The $\alpha^{-2}$ parts of (transformed) eqs. (36) lead to equations on $\overline{\lambda}_{a}$ with the solution
$\overline{\lambda}_{1}=-2,\quad\overline{\lambda}_{2}=-2,\quad\overline{\lambda}_{3}=2,$ (40)
while the $\alpha^{-1}$ and $\alpha^{0}$ parts of the equations $\{0=R_{11},0=R_{12},0=R_{21}\}$ (after substituting eqs. (40) and multiplying by $\alpha$) lead to recursive relations
$\overline{\bf I}_{1}^{+}g(\underline{n})$ $=$ $\big(-2\overline{\bf I}_{1}^{-}-6\overline{\bf I}_{2}^{-}+2\overline{\bf I}_{3}^{-}+\alpha\,(3\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}+2\overline{\bf I}_{1}^{-}\overline{\bf I}_{3}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{1}^{+}-2\overline{\bf I}_{2}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{1}^{+})\big)g(\underline{n}),$
$\overline{\bf I}_{2}^{+}g(\underline{n})$ $=$ $\big(-6\overline{\bf I}_{1}^{-}-2\overline{\bf I}_{2}^{-}+2\overline{\bf I}_{3}^{-}+\alpha\,(2\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}+\overline{\bf I}_{1}^{-}\overline{\bf I}_{2}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{2}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{2}^{+})\big)g(\underline{n}),$
$\overline{\bf I}_{3}^{+}g(\underline{n})$ $=$ $\big(2\overline{\bf I}_{1}^{-}+2\overline{\bf I}_{2}^{-}-2\overline{\bf I}_{3}^{-}+\alpha\,(-2\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}-\overline{\bf I}_{1}^{-}\overline{\bf I}_{3}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{3}^{+})\big)g(\underline{n}).$ (41)
Eqs. (38-41) allow us to derive a relation between $f(\underline{n})$ for a particular $(\underline{n})$ and $f(1,1,1)$; as an example, let us do this for $f(2,1,1)$. In the first step we need to express $f(2,1,1)$ and $f(1,1,1)$ in terms of $g(\underline{n})$. According to eq. (38) with (40)
$f(2,1,1)$ $=$ $(1-2\alpha^{-2}\overline{\bf I}_{1}^{-}+\dots)f^{\prime}(2,1,1)$ (42)
$=$ $f^{\prime}(2,1,1)-2\alpha^{-2}f^{\prime}(1,1,1)$
$=$ $\alpha^{-1}g(2,1,1)-2\alpha^{-2}g(1,1,1),$
$f(1,1,1)$ $=$ $g(1,1,1).$ (43)
From here on, dots denote terms that do not contribute due to the conditions (37, 39). In the second step we reduce the right-hand side of eq. (42) to $g(1,1,1)$. The first equation (41) for $(\underline{n})=(1,1,1)$ gives
$g(2,1,1)=(3\alpha\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}+\dots)g(1,1,1)=3\alpha g(1,1,1),$ (44)
and hence
$f(2,1,1)=(3-2\alpha^{-2})g(1,1,1)=(3-d)g(1,1,1).$ (45)
From eqs. (43, 45) we finally obtain
f(2,1,1)=(3d)f(1,1,1).f(2,1,1)=(3-d)f(1,1,1). (46)
Please note that in this example the recursion stopped due to condition (39), so the result (in terms of $\alpha$) is exact; in general one should stop when the desired order of $\alpha$ is reached (for compactness we do not expand $g(\underline{n})$ into a series in $\alpha$ explicitly). Also note that only even powers of $\alpha$ (integer powers of $d$) contribute to the final relation (46), unlike the relations (42, 44).
The standard Laporta’s approach, that is solving eqs. (36, 37) for some (sufficiently large) set of values for $(\underline{n})$, gives
$0$ $=$ $R_{11}(1,1,1)-R_{21}(1,1,1)$
$=$ $1/2(d-2\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}-\overline{\bf I}_{2}^{-}\overline{\bf I}_{2}^{+}+\overline{\bf I}_{2}^{+}+\dots)f(1,1,1)$
$=$ $(d/2-3/2)f(1,1,1)+1/2f(2,1,1),$
which confirms eq. (46).
@@
Figure 1 presents “hard” $\overline{\lambda}_{a}$ (“soft” $\tilde{\lambda}_{a}$ can be recovered using eqs. (32)) for leading sectors of some integral families up to the 7-loop order. Polynomials of the first degree as coefficients were sufficient to reduce the original system of equations to an equation for one variable. [^2] Symbolic manipulations were performed using REDUCE FreeREDUCE ).
L=5L=5cceeeecceeeebbbbaabbbbcc
L=6L=6cceeeeeecceeeeeebbbbaabbbbcccc
L=7L=7cceeeeeeeecceeeeeeeebbbbaabbbbcccccc
L=6L=6eeeeeeeeeeeeccccccccbbbbaabbbb
L=7L=7eeeeeeeeeeeeeeeeccccccccccbbbbaabbbb
Figure 1: $\overline{\lambda}_{a}$ (roots of eqs. (24)) for some vacuum diagrams. Ordinary lines are massless, double line is massive with $\mu=1$; $a_{L}=L$, $b_{L}=-\frac{L-2}{2}e_{L}$, $c_{L}=\frac{L(L-2)}{2}e_{L}-L(L-1)$;
@@
The results show that in the 5-, 6-loop planar examples there is at least one variant of the substitutions which transform IBP relations to a recursive form, in the 7-loop planar and in the 6-, 7-loop nonplanar cases there are at least two. The $L$-dependence of the results was obtained empirically, but there are some indications of its validity for an arbitrary order, we hope to clarify this point in future publications.
Please note that eqs. (16) from Baikov:2005nv (eqs. (34) in the notation of this article) at the 7-loop level lead to polynomial equations of the 6th degree for 10 variables, which is a much more difficult task.
@@ 6 Conclusion
In this paper we present a method to construct solutions of IBP relations for Feynman integrals as formal series in the limit $1/d\rightarrow 0$. The idea is to use substitutions of the form (1517) which transform IBP relations into an explicitly recursive form. From a technical point of view, it is necessary to calculate the constants $\lambda$ as solutions to a system of bilinear quadratic equations (24). Although this is relatively simple (and therefore can be automated) in the case where one line is massive and the others are massless, more general kinematic configurations may be more difficult to study. In particular, it may happen that the number of $1/d$ solutions constructed will be less than the number of master integrals and, therefore, not enough to calculate all $c_{i}(\underline{n})$ coefficients in eq. (7). However, even in this case they provide some linear relations between $c_{i}(\underline{n})$ that can be used (in combination with other data) for their reconstruction or at least for verification.
Please note that although in this article we are dealing with recurrence relations for vacuum integrals, the substitutions (15, 17) can also be applied to the case with external momenta. In this context, it is worth mentioning the work Mizera:2019vvs , where the large $d$ limit was used to obtain the differential equations for master integrals. We hope that studying possible transformations of the corresponding IBP relations will help in the difficult task of evaluating “multi-legs” integrals; we leave this for future research.